OpenStudy (anonymous):

find three concecutive even integers such that 3 times the sum of the first two is 48 less than the third

OpenStudy (callisto):

Let x be the smallest integer required, the next two would be (x+2), (x+4) Since 3 times the sum of the first two is 48 less than the third. 3[ x + (x+2) ] =( x+4 ) - 48 3( 2x +2) = x - 44 6x + 6 = x- 44 Can you solve it from here?

OpenStudy (anonymous):

im sorry im still a bit confused

OpenStudy (callisto):

Which step?

OpenStudy (anonymous):

ok what is the equasion for the first step?

OpenStudy (callisto):

3(3) times (x) sum of the first two number [ x + (x+2) ] is (=) 48 less than (-48) the third (x+4) So, you can write it into an equation as 3 x [ x + (x+2)] = (x+4)-48 ^ ( times, not unknown x) Got this step?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

And do you understand the rest of the steps?

OpenStudy (anonymous):

i think so

OpenStudy (callisto):

That's good :) Can you work out the answer now?

OpenStudy (anonymous):

i let u know in a few min.:)

OpenStudy (callisto):

Okay :)

OpenStudy (anonymous):

ok so i got 4x=-46:(

OpenStudy (callisto):

o.o .... 3[ x + (x+2) ] =( x+4 ) - 48 3( 2x +2) = x - 44 6x + 6 = x- 44 6x -x +6 = x-x -44 5x +6 = -44 5x +6-6 = -44-6 5x = -50 ....